OpenStudy (1davey29):

Let f be the function defined by f(x)=ln(3x+2)^k for some positive constant k. If f'(2)=3, what is the value of k?

OpenStudy (1davey29):

(A) ln3/ln8 (B) ln8 (C) 4 (D) 8 (E) 16

OpenStudy (phi):

do you know how to find the derivative of \[ \ln (3x+2)^k \] ?

OpenStudy (1davey29):

I believe so. Is it this? \[k(\ln(3x+2))^{k-1}*\frac {3}{3x+2}\]

OpenStudy (phi):

almost. d (ln u) = 1/u du in this case u = (3x+2)^k 1/u is 1/(3x+2)^k what is du ?

OpenStudy (phi):

in other words, what is \[ \frac{d}{dx} (3x+2)^k \]?

OpenStudy (1davey29):

k(3x+2)^(k-1). So the derivative of ln(3x+2)^k is (k(3x+2)^(k-1))/(3x+2)?

OpenStudy (phi):

still missing it. You should have a 1/u factor where u is (3x+2)^k \[ \frac{1}{(3x+2)^k }\cdot k(3x+2)^{k-1} \cdot 3 \]

OpenStudy (phi):

now simplify that, and plug in x=2

OpenStudy (1davey29):

Ok. So it would end up as this? \[\frac{ 1 }{ (3(2)+2)^k }*k(3(2)+2)^{k-1}*3=\frac{3k(8)^{k-1}}{8^k}\]

OpenStudy (phi):

ok, but it's nicer to write (3x+2)^k-1 / (3x+2)^k as 1/(3x+2) first but using what you did, notice that 1/8^k is 8^(-k) and 8^k * 8^-1 * 8^-k = 8^-1 or 1/8 so you have 3k/8 and they say that f'(2)=3 so 3k/8 = 3

OpenStudy (1davey29):

I see. So now I just solve for k, which is 8, correct?

OpenStudy (phi):

yes

OpenStudy (1davey29):

Ok. Thanks for the help.

OpenStudy (phi):

so this question was a bit complicated. First, you need to know how to take the derivative of ln and use the chain rule. That can be confusing without practice. 2nd, you should be comfortable with knowing how to simplify expressions with exponents For example, my mind immediately simplified the (3x+2)^k-1 / (3x+2)^k i.e. look for those patterns and simplify first.